Green Airways flies several daily flights from Dallas to Chicago with 150-seat planes. Based on historical data, the flight on Wednesday evening before Thanksgiving is always sold-out. However, there are usually no-shows, so the airline decides to improve revenue by overbooking. The distribution of no-shows is according to the table below and the airline estimates that the cost of bumping a passenger is about 10 times more than the ticket price. How many seats should Green Airways overbook? Please show work. Thank you.

Number of no-shows

Probability

0

0.01

1

0.10

2

0.63

3

0.26
Answer #1

Let the ticket price = P

So, the cost of bumping = 10*P

Cost of overbooking one less seat than the number of no-shows, Cu = P
Cost of overbooking one more seat than the number of no-shows, Co = 10P

So, the optimal in-stock probability = Cu / (Co + Cu) = P / (10P + P) = 0.09

Number of no-shows

Probability

Cumulative probability

0

0.01

0.01 < 0.09

1

0.10

0.11 > 0.09

2

0.63

0.74

3

0.26

1.00

The cumulative probability crosses the optimal in-stock probability for the number of no-shows=1

So, the optimal overlooking = 1